OpenStudy (anonymous):

Help? The point P(18, 24) is on the terminal side of θ. Evaluate sin θ.

OpenStudy (anonymous):

4/5

OpenStudy (anonymous):

thats if you didn't forget the hypotenuse

OpenStudy (anonymous):

Can you explain how you did it so I will know for future reference? c:

OpenStudy (anonymous):

r^2=18^2+24^2 r^2=900 r=30

OpenStudy (anonymous):

i can't really explain the rest without the equation editor. it isn't working. sorry c;

OpenStudy (anonymous):

It's fine. Thanks anyways! God bless you! :D

OpenStudy (anonymous):

:)